# Facing Asymmetry - Uncovering the Causal Link between Facial Symmetry and Expression Classifiers using Synthetic Interventions

### v1

@@ 1 Introduction
Emotional expressiveness is a crucial topic in our daily life for communicating our internal state and for understanding other people [95, 76]. The state-of-the-art for automatically classifying the six base emotions [23] is achieved by end-to-end trained black box neural networks [65, 82, 110, 102, 24, 17]. However, it remains unclear how the internal decision-making processes of these models respond to out-of-distribution inputs that were not accounted for during training. Specifically, we observe a performance degradation when classifying facial expressions in individuals with unilateral facial palsy, a condition impairing the ability to produce symmetrical facial expressions due to underlying nerve damage. Although intuition suggests that facial asymmetry could influence model behavior, we lack a quantifiable way to test and validate this hypothesis.
@@ 3.1 Preliminaries & Causal Point of View
Causal inference tries to answer causal questions from data [64]. This includes interventions, i.e., additional experiments, and purely observational data. Importantly, causal questions can be categorized into a hierarchy. This so-called Pearl’s causal hierarchy (PCH) [4] consists of the following three levels ordered by increasing difficulty: associational, interventional, and counterfactual questions. The latter two are analyzed using the do-operator [60], which changes a variable to a constant value, e.g., we would write $do(\text{Facial Symmetry}:=s)$ for the variable facial symmetry.
@@ 4 Experiments and Results
Compared to other research that requires training [43, 15], we can asses 17 existing expression classifiers. Please note that our selection of models is not intended to be comprehensive or representative of all possible classifiers but as a first case study. We focus on a subset of models that provide code and model weights, which are likely to be applied more frequently in out-of-domain scenarios like medicine and psychology. To test our hypothesis, we perform two groups of experiments. First, we investigate if a symmetry bias exists for real-world data inside classifiers using associational methods [73, 71, 61, 7]. Second, moving up on the causal hierarchy [4], we connect facial symmetry and model output using our synthetic intervention framework and score.
@@ 4.1 Experiment 1: Observations on Real-World Data
The participants’ expressions are captured while mimicking the six basic emotions four times in a random order [32, 6, 9, 7] following FER2013 [20]. We focus on the happy expression as it most impacts emotional expressiveness [76, 95, 8]. Examples of each face group are shown in Fig. 4. We measure an average classification accuracy of $97.40\%$ for probands and $58.25\%$ for patients among the FER2013 models (due to the similar recording setup), indicating the decreased performance under facial palsy. A more fine-grained analysis can be found in the supplementary material.
@@
Table 1: Significance results ($p<0.01\rightarrow$ ) of [73] on our data for three symmetry features. We investigate happy logits with respect to the binary facial palsy state and LPIPS [108] similarity between the face sides for real-world images of facial palsy patients and healthy probands.
Following related work [9, 73, 61], we denote all significant behavior changes in Table 1 using the majority decision of three conditional independence tests [27, 79, 11] (for detailed hyperparameters see supplementary material). We find all 17 models show a statistically significant shift in their happy activations for varying facial symmetries in the real-world data. Please note that DAN [102], trained on RAFDB [47, 46], is the only classifier where we find no significance regarding binary facial palsy, which is a highly discretized form of symmetry. However, the continuous symmetry measure LPIPS [108] indicates the same behavior changes. These results provide first evidence for our hypothesis that expression classifiers are biased toward facial symmetry, especially concerning downstream applications.
Furthermore, we provide more qualitative investigations of the output changes in the supplementary material. For most classifiers, we observe, on average, a decrease in activations. Given the performance decrease, this observation is expected and indicates uncertainty in predicting the happy class. However, these are associational investigations [4], i.e., we do not isolate changes in facial symmetry from confounding factors. Hence, while we observe changes in classifier behavior on real-world data, our following interventional investigation in Section 4.2 provides more reliable, actionable insights.
@@ 4.2 Experiment 2: Synthetic Facial Symmetry Interventions
First, we observe that facial symmetry impacts each expression’s logit activation for the models irrespective of the dataset. The ResidualMaskinNet [65] trained on FER2013 [20] does not reach high activations for fear and sad. Further, they decrease even more for lower symmetry values. Expressions such as angry, disgust, or surprise have higher activations and seem to be affected only by more pronounced asymmetry. Especially fear also seems problematic for the other classifiers highlighted in Fig. 5. Notably, the same behavior holds for HSEmotion [82] and DAN [102]: Lower symmetry leads to lower output activations for all expressions, providing further evidence to confirm our hypothesis.
@@
Next, we go one step further and quantitatively measure the impact of facial symmetry using our proposed score (Section 3.3). Table 2 summarizes these results averaged over all individuals $\mathfrak{I}$. Remember that a positive score corresponds to increased activations for increased symmetry. Further, we report the ratio of significant systematic changes over the set of individuals $\mathfrak{I}$ in Table 1 of the supplementary material. In most cases, we observe a significant impact of facial symmetry for all classifiers and expressions. Hence, this enables us to interpret the patterns we observe in Table 2 concerning the expressions and training datasets. We note the highest impact of $0.0373$ for surprise of PosterV2 [50]. This score indicates that, on average, over the complete surprise onset, increasing the symmetry by one step in our simulation increased the softmax output of PosterV2 by $3.7$ percentage points. However, while all 17 classifiers are significantly impacted by changes in facial symmetry, the effect size can still be small (see, for example, fear in Figure 5(c)).
@@ 5 Limitations and Social Impact
Our work relies on the statistical shape model, which induces a domain shift. While many classifiers advertise the out-of-domain usage, e.g., [84, 83, 82, 65], and we additionally optimize the synthetic faces regarding model and expression (see Section 3.2). Nevertheless, other factors like camera angle could jointly influence the behavior. Secondly, while we focus on facial symmetry, other facial features, e.g., age or skin color, can impact classifier performance and should be considered in the future. Our framework cannot model the entire range of facial nerve damage effects on facial mimicry, as we do not account for all possible forms of facial asymmetry, e.g., synkinetic effects.
Regarding societal impact, our work investigates only existing expression classifiers. We move from the associational level to causal interventions to better understand how these black-box models operate. Such an approach could benefit other disciplines, primarily psychological and medical applications. We provide our experiments’ generative framework and evaluation code so that other researchers can simply evaluate their models.
@@ 6 Conclusion
We tested our hypothesis on real-world data of facial palsy patients and healthy probands using associational methods [73, 61, 9]. Here, we saw that facial palsy and the similarity between the face halves led to 33 out of 34 tests being significant. To verify these results in a more controlled manner, moving up the causal hierarchy, we employed our interventional framework to test the impact of symmetry on the models’ behavior. We observed that many classifiers, on average, decrease logit activations for lower facial symmetry.
@@ 6.0.1 Ethics Approval
Written informed consent was obtained from all participants, including displayed individuals. The ethics committee of the Jena University Hospital approved the study (No. 2019-1539).
@@ 6.0.2 Acknowledgement
6.0.3 Disclosure of Interests.
The authors have no competing interests to declare.
@@ Appendix 0.A Structural Causal Models
A structural causal model (SCM) is defined as a 4-tuple $M=(U,V,F,P)$, where $U$ is a set of exogenous variables describing outside factors, $V=\{V_{1},...,V_{n}\}$ is the set of endogenous variables we measure in our model, $F=\{f_{1},...,f_{n}\}$ is a set containing functions $f_{i}$ that describe the functional relationships, and $P$ is a joint probability distribution over $U$. Further, each $V_{i}$ has a set of parents $PA_{i}$ that functionally determine $V_{i}$ together with some exogenous variables $U_{i}\subseteq U$. These parents $PA_{i}$ are a subset of $V\setminus\{V_{i}\}$. For settings $pa_{i}$ of parents $PA_{i}$ and $u_{i}$ of the exogenous variables $U_{i}$, $f_{i}$ determines the value $v_{i}=f_{i}(pa_{i},u_{i})$ of $V_{i}$.

### v2

@@ 1 Introduction
Emotional expressiveness is a crucial topic in our daily life for communicating our internal state and for understanding other people [95, 76]. The state-of-the-art for automatically classifying the six base emotions [23] is achieved by end-to-end trained black box neural networks [65, 82, 110, 102, 24, 17]. However, it remains unclear how the internal decision-making processes of these models respond to out-of-distribution inputs due to likely unbalanced training data. Specifically, we observe a performance degradation when classifying facial expressions in individuals with unilateral facial palsy, a condition impairing the ability to produce symmetrical facial expressions due to underlying nerve damage. Although intuition suggests that facial asymmetry could influence model behavior, we lack a quantifiable way to test and validate this hypothesis.
@@ 3.1 Preliminaries & Causal Point of View
Causal inference tries to answer causal questions from data [64]. This includes interventions, i.e., additional experiments and purely observational data. Importantly, causal questions can be categorized into a hierarchy. This so-called Pearl’s causal hierarchy (PCH) [4] consists of the three levels ordered by increasing difficulty: associational, interventional, and counterfactual questions. The latter two are analyzed using the do-operator [60], which changes a variable to a constant value, e.g., we write $do(\text{Facial Symmetry}:=s)$ for the variable facial symmetry.
@@ 4 Experiments and Results
We can assess existing expression classifiers, whereas other research requires training [43, 15]. Please note that our selection of models is not intended to be comprehensive or representative of all possible classifiers but as a first case study. We focus on a subset of models that provide code and model weights, which are likely to be applied in out-of-domain scenarios like medicine and psychology. To test our hypothesis, we perform two groups of experiments. First, we investigate if a symmetry bias exists for real-world data inside classifiers using associational methods [73, 71, 61, 7]. Second, moving up on the causal hierarchy [4], we link facial symmetry and model output using our synthetic intervention framework.
@@ 4.1 Experiment 1: Observations on Real-World Data
The participants’ expressions are captured while mimicking the six basic emotions four times in a random order [32, 6, 9, 7] following FER2013 [20]. We focus on the happy expression as it most impacts emotional expressiveness [76, 95, 8]. Examples of each face group are shown in Fig. 4. We measure an average classification accuracy of $97.40\%$ for probands and $58.25\%$ for patients among the FER2013 models, indicating the decreased performance under facial palsy. A more detailed analysis can be found in the supplementary material.
@@
Table 1: Significance results ($p<0.01\rightarrow$ ) of [73] on our data for three symmetry features. We analyze happy logits regarding the binary facial palsy state and LPIPS [108] similarity between the face sides for facial palsy patients and healthy probands images.
Following related work [9, 73, 61], we denote all significant behavior changes in Table 1 using the majority decision of three conditional independence tests [27, 79, 11] (for detailed hyperparameters see supplementary material). We find all 17 models show a statistically significant shift in their happy activations for varying facial symmetries in the real-world data. Please note that DAN [102], trained on RAFDB [47, 46], is the only classifier where we find no significance regarding binary facial palsy, which is a highly discretized form of symmetry. However, the continuous symmetry measure LPIPS [108] indicates the same behavior changes. These results provide evidence for our hypothesis that expression classifiers are biased toward facial symmetry, especially concerning downstream applications.
We provide more qualitative investigations of the output changes in the supplementary material. For most classifiers, we observe, on average, a decrease in activations. Given this decrease, this observation is expected and indicates uncertainty in predicting the happy class. However, these are associational investigations [4], i.e., we cannot isolate changes due to only facial symmetry. Hence, while we observe changes in classifier behavior on real-world data, our interventional investigation in Section 4.2 provides more reliable, actionable insights.
@@ 4.2 Experiment 2: Synthetic Facial Symmetry Interventions
We observe that facial symmetry impacts each expression’s logit activation for the models irrespective of the dataset. The ResidualMaskinNet [65] trained on FER2013 [20] does not reach high activations for fear and sad. Further, they decrease even more for lower symmetry values. Expressions such as angry, disgust, or surprise have higher activations and seem to be affected only by more pronounced asymmetry. Especially fear also seems problematic for the other classifiers highlighted in Fig. 5. Notably, the same behavior holds for HSEmotion [82] and DAN [102]: Lower symmetry leads to lower output activations for all expressions, providing further evidence to confirm our hypothesis.
@@
We go one step further and quantitatively measure the impact of facial symmetry using our proposed score (Section 3.3). Table 2 summarizes these results averaged over all individuals $\mathfrak{I}$. A positive score corresponds to increased activations for increased symmetry. Further, we report the ratio of significant systematic changes over the set of individuals $\mathfrak{I}$ in Table 1 of the supplementary material. In most cases, we observe a significant impact of facial symmetry for all classifiers and expressions. This enables us to interpret the patterns we observe in Table 2 concerning the expressions and training datasets. We note the highest impact of $0.0373$ for surprise of PosterV2 [50]. This score indicates that, on average, over the complete surprise onset, increasing the symmetry by one step in our simulation increased the softmax output of PosterV2 by $3.7$ percentage points. However, while all 17 classifiers are significantly impacted by changes in facial symmetry, the effect size can still be small (see, for example, fear in Figure 5(c)).
@@ 5 Limitations and Social Impact
Our work relies on the statistical shape model, inducing a domain shift and limiting possible expressions. Some classifiers advertise the out-of-domain usage, e.g., [84, 83, 82, 65], and we optimize the synthetic faces regarding model and expression (see Section 3.2). Factors like camera angle could jointly influence the behavior. Secondly, other facial features, e.g., age or skin color, could impact classifier performance and should be considered. In our current framework, we cannot account for all possible forms of facial asymmetry, e.g., synkinetic effects.
Regarding societal impact, we investigate existing expression classifiers only. We move from the associational level to causal interventions to better understand how these black-box models operate. This could benefit other disciplines, primarily psychological and medical applications. We provide our experiments’ framework and evaluation code so that researchers can evaluate their models.
@@ 6 Conclusion
We tested our hypothesis on real-world data using associational methods [73, 61, 9]. Here, we saw that facial palsy and the similarity between the face halves led to 33 out of 34 tests being significant. To verify these results in a controlled manner, moving up the causal hierarchy, we employed our interventional framework to test the impact of symmetry on the models’ behavior. We observed that many classifiers, on average, decrease logit activations for lower facial symmetry.
@@ 6.0.1 Ethics Approval
Written consent was obtained from all participants. The Jena University Hospital ethics committee approved the study (No. 2019-1539).
@@ Appendix 0.A Structural Causal Models
A structural causal model (SCM) is defined as a 4-tuple $M=(U,V,F,P)$, where $U$ is a set of exogenous variables describing outside factors, $V=\{V_{1},...,V_{n}\}$ is the set of endogenous variables we measure in our model, $F=\{f_{2},...,f_{n}\}$ is a set containing functions $f_{i}$ that describe the functional relationships, and $P$ is a joint probability distribution over $U$. Further, each $V_{i}$ has a set of parents $PA_{i}$ that functionally determine $V_{i}$ together with some exogenous variables $U_{i}\subseteq U$. These parents $PA_{i}$ are a subset of $V\setminus\{V_{i}\}$. For settings $pa_{i}$ of parents $PA_{i}$ and $u_{i}$ of the exogenous variables $U_{i}$, $f_{i}$ determines the value $v_{i}=f_{i}(pa_{i},u_{i})$ of $V_{i}$.